OpenStudy (anonymous):

$1,600 principal earning 7%, compounded semi-annually, after 33 years. Someone really explain and go through this with me.

OpenStudy (wolf1728):

First we should determine what 7% semi-annual equals annual = (1+.07/2) ^2 = 1.071225 annual interest = 7.1225%

OpenStudy (wolf1728):

Basically, it depends on using the formulas correctly.

OpenStudy (anonymous):

Okay, thank you :)

OpenStudy (wolf1728):

total = principal*(1+rate)^33 total = 1,600 * (1.071225)^33 total = 1,600 * 9.6841852012 total = 15,494.70 (Double check your work with this calculator: http://1728.org/compint.htm